# How Sparse Can We Prune A Deep Network: A Fundamental Limit Perspective

### v3

@@ D.2 Gaussian Width of the Deformed Ellipsoid
′′$\mathcal{L}^{{}^{\prime\prime}}({\bf w})$ =′′$=\mathcal{L}^{{}^{\prime\prime}}({\bf w}_{0}+{\bf v})$
@@
where ′′$\mathcal{L}^{{}^{\prime\prime}}({\bf w})=\hat{\bf H}$, Let ${\bf H}=\hat{\bf H}+{\bf H}_{0}$ with ${\bf H}_{0}$ is closed to ${\bf 0}$, considering the Weyl inequality:
@@
where $\|{\bf H}_{0}\|_{2}$ is small enough. So if $\hat{\lambda}_{i}(\hat{{\bf H}})$ is less than 0, since $\hat{\lambda}_{i}(\hat{{\bf H}})\geq\lambda_{i}({\bf H})-\|{\bf H}_{0}\|_{2}$, its absolute value $|\hat{\lambda}_{i}(\hat{{\bf H}})|\leq\|{\bf H}_{0}\|_{2}-\lambda_{i}({\bf H})\leq\|{\bf H}_{0}\|_{2}$, which means that the negative eigenvalues of the Hessian matrix have small magnitudes.
@@
Consider a well-trained neural network with weights ${\bf w}$, whose loss function defined by ${\bf w}$ has a Hessian matrix ${\bf H}$. Due to the non-positive definiteness of ${\bf H}$, there exist negative eigenvalues. Let the eigenvalue decomposition of ${\bf H}$ be ${\bf H}={\bf v}^{T}{\bf\Sigma v}$, where ${\bf\Sigma}$ is a diagonal matrix of eigenvalues. Let ${\bf D}={\bf v}^{T}{\bf|\Sigma|v}$, where ||$\cdot$|| ⋅ | means absolute operation. the geometric objects defined by H and D are $S(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf H}{\bf w}\leq\epsilon\}$ and $\hat{S}(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf D}{\bf w}\leq\epsilon\}$, then the gaussian width of the two set satisfy:
@@ Appendix E Comparison between the Upper and Lower Bound
E.1 $D-k$ Dimension Sublevel Set is Still an Ellipsoid
In the derivation of the upper bound for the pruning ratio threshold, we employed a $D-k$ dimensional loss sublevel set:
@@ Appendix F Sharpness & Lower Bound
[…]
@@ G.2 Comparison of Pruning Algorithms
CIFAR10 FC5 55.3$\pm$0.62 1.7 59.96$\pm$0.45 58.76$\pm$0.15 58.83$\pm$0.24 -
FC12 55.5$\pm$0.26 1.0 60.84$\pm$0.21 54.96$\pm$0.28 59.37$\pm$0.21 -
VGG16 90.73$\pm$0.22 0.6 91.66$\pm$0.08 88.76$\pm$0.13 89.22$\pm$0.24 87.32$\pm$0.11
CIFAR100 ResNet18 72.19$\pm$0.23 1.9 71.82$\pm$0.09 69.33$\pm$0.22 68.55$\pm$0.21 65.74$\pm$0.27
ResNet50 74.07$\pm$0.43 2.0 75.22$\pm$0.11 72.21$\pm$0.25 69.02$\pm$0.23 68.58$\pm$0.33
TinyImageNet ResNet18 52.92$\pm$0.13 4.2 55.42$\pm$0.02 45.43$\pm$0.27 45.12$\pm$0.19 43.28$\pm$0.21
ResNet50 56.45$\pm$0.17 2.9 57.49$\pm$0.01 51.41$\pm$0.28 46.93$\pm$0.41 40.42$\pm$0.31

### v5

@@ D.2 Gaussian Width of the Deformed Ellipsoid
$\mathcal{L}^{{}^{\prime\prime}}({\bf w})$ $=\mathcal{L}^{{}^{\prime\prime}}({\bf w}_{0}+{\bf v})$
@@
where $\mathcal{L}^{{}^{\prime\prime}}({\bf w})=\hat{\bf H}$, Let ${\bf H}=\hat{\bf H}+{\bf H}_{0}$ with ${\bf H}_{0}$ is closed to $\bf 0$}, considering the Weyl inequality:
@@
[…]
@@
Consider a well-trained neural network with weights $\bf w$}, whose loss function defined by $\bf w$} has a Hessian matrix $\bf H$}. Due to the non-positive definiteness of $\bf H$}, there exist negative eigenvalues. Let the eigenvalue decomposition of $\bf H$} be ${\bf H}={\bf v}^{T}{\bf\Sigma v}$, where $\bf\Sigma$} is a diagonal matrix of eigenvalues. Let ${\bf D}={\bf v}^{T}{\bf|\Sigma|v}$, where $|\cdot|$ means absolute operation. the geometric objects defined by H and D are $S(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf H}{\bf w}\leq\epsilon\}$ and $\hat{S}(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf D}{\bf w}\leq\epsilon\}$, then the gaussian width of the two set satisfy:
@@ Appendix E Comparison between the Upper and Lower Bound
E.1 DkD-k Dimension Sublevel Set is Still an Ellipsoid
In the derivation of the upper bound for the pruning ratio threshold, we employed a DkD-k dimensional loss sublevel set:
@@ Appendix F Sharpness & Lower Bound
$\leq 1-\frac{2\epsilon}{D}\frac{D^{2}}{\sum_{i=1}^{D}(\|{\bf w}^{*}-{\bf w}^{k}\|_{2}^{2}\lambda_{i}+2\epsilon)}\quad\quad\quad\quad\text{Cauchy–Schwarz Inquality.}$
@@ G.2 Comparison of Pruning Algorithms
[…]